For the next two questions, consider Adulttown, where everyone is of working age. It has a population of 50,000. Of those 50,000:

38,000 work full-time
6000 work part-time
4500 aren't working or looking for work
1500 aren't working but are looking actively for work.

Calculate this economy's labor force participation rate, carefully following all numeric instructions. Enter your answer "as a percent, but without the percentage sign." In other words, if your answer is 0.9999 (99.99 percent), enter just 99.99 in the blank.
Answer #1

Labor force = employed+unemployed

= 38000+6000+1500

= 45500

Adult population = 50000

LFPR = 45500/50000*100= 91%